OpenStudy (anonymous):

2/3-x/9=2 Solve for x

OpenStudy (ddcamp):

2/3 = 6/9, so we have: (6/9) - (x/9) = 2 (6-x)/9 = 2 6-x = 18 -x = 12 x = -12

OpenStudy (anonymous):

thank you !!!